OpenStudy (anonymous):

Cindy is measuring the area of several parallelograms. Which two parallelograms have the same area?

OpenStudy (anonymous):

I don't see any parallelograms?

OpenStudy (anonymous):

im tryin to post it

OpenStudy (dan815):

the area of a parallelograms is higher * base so which every paralleograms have the same height and base are equal

OpenStudy (dan815):

heighet * base

OpenStudy (anonymous):

doesnt let me post da images how can we post it on here?

jimthompson5910 (jim_thompson5910):

you click on "attach file" to post the images

OpenStudy (anonymous):

i did but da images dont show on my computer how can we save it to our computer?

jimthompson5910 (jim_thompson5910):

you can take a screenshot

jimthompson5910 (jim_thompson5910):

do you know how to do that?

OpenStudy (dan815):

jim thompson can you help me with some math plz

OpenStudy (anonymous):

lmaooo im confused...how do we screen shot?

jimthompson5910 (jim_thompson5910):

you hit the printscreen key then you open up MS paint and paste in the image save it, then send it

OpenStudy (anonymous):

ok hold on

jimthompson5910 (jim_thompson5910):

ok

jimthompson5910 (jim_thompson5910):

ok great, thanks

jimthompson5910 (jim_thompson5910):

do you know how to find the area of a parallelogram?

OpenStudy (anonymous):

there it is

jimthompson5910 (jim_thompson5910):

yes? or no?

OpenStudy (anonymous):

sorry the system went down for a moment

jimthompson5910 (jim_thompson5910):

yeah it does that to me a lot

jimthompson5910 (jim_thompson5910):

were you able to find out the area of each?

OpenStudy (anonymous):

i kno how to find a perimeter but not an area

jimthompson5910 (jim_thompson5910):

Area = Base * Height

jimthompson5910 (jim_thompson5910):

so the first one on the very left for example is A = b*h A = 14*16 A = 224

OpenStudy (anonymous):

the base is the number outside or inside the figure?

jimthompson5910 (jim_thompson5910):

the base is either the bottom or the top length

OpenStudy (anonymous):

oooo ok

jimthompson5910 (jim_thompson5910):

since bottom = top in any parallelogram

OpenStudy (anonymous):

ok let me find the area of each real quick

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

1.224 2.63 3.63 4.182 soooo 2 and 3 are equal figures

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

thank u!!!!!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

i did the other ones how u told me..and i got dem wrong...?????im confused now.. i got 3 of dem dat i did by myself right and da others wrong?

jimthompson5910 (jim_thompson5910):

show me which ones you got wrong

OpenStudy (anonymous):

The figure below shows the shape of a boat’s sail. What is the area of the sail?

jimthompson5910 (jim_thompson5910):

ok this is a triangle, so you use a different formula this time

jimthompson5910 (jim_thompson5910):

A = b*h/2 b = base h = height

jimthompson5910 (jim_thompson5910):

that's the formula you use this time

OpenStudy (anonymous):

:((((((

jimthompson5910 (jim_thompson5910):

it's not too different from the one before

jimthompson5910 (jim_thompson5910):

tell me what you get when you use that new formula